OpenStudy (anonymous):

quick computing company produces calculators they have found that the cost c(x) of making x calculators is a quadratic function in terms of x. The company also discovered that it costs $45.00 to produce two calculators $143 to produce four calculators $143 to produce four calculators and $869 to produce ten calculators. Find the cost of producing seven calculators

OpenStudy (anonymous):

Hey old school mate! I know you

OpenStudy (anonymous):

You look familiar

OpenStudy (anonymous):

lol that was a prank to get you started with a bit of funXD

OpenStudy (anonymous):

lol you look like a guy in my class

OpenStudy (anonymous):

In question like this you need to start organizing the known values and keys that can lead to finding the final answer.

OpenStudy (anonymous):

First you know that this is a quadratic equation we all know to be old school nostalgia

OpenStudy (anonymous):

Ok since now that you know this is a quadratic function you find the coordinates to find in with other coordinates in the form of quadratic equation which has parabola and vertex just like your teacher's lesson

OpenStudy (anonymous):

you need to assume that this is a quadratic equation and that all of the coordinates described conform with the shape of that parabola old school

OpenStudy (anonymous):

You can either use your calculator to do it, just assign it to quadratic equation and enter the coordinates given. Or you can draw it on a cartesian plane to get the approximate amount.

OpenStudy (anonymous):

Or you can go about calculating it by hand which would take a bit of trick ;) Let me know how it goes.

OpenStudy (anonymous):

um.. what?

OpenStudy (anonymous):

Good luck:)

OpenStudy (anonymous):

math stands for mental abuse to humans;) and you are the participant;)

OpenStudy (anonymous):

I have no clue how to do what you described above

OpenStudy (anonymous):

Ok. Which part? I don't understand.

OpenStudy (anonymous):

any of it

OpenStudy (welshfella):

let c(x) = ax^2 + bx + c general form of a quadratic function from the data 45 = a(2)^2 + 2b + c 143 = a(4)^2+ 4b + c 869 = a(10^2 + 10b + c from this we can find values of a , b and c then plug in x = 7

OpenStudy (welshfella):

so the system of equations is 4a + 2b + c = 45 16a + 4b + c = 143 100a + 10b + c = 869

OpenStudy (welshfella):

looks a bit scary but they can be solved

OpenStudy (welshfella):

if you subtract the first one from the second then 2nd from the third you will eliminate the c and get 2 equations in a and b

OpenStudy (welshfella):

12a + 2b = 98 84a + 6b = 726

OpenStudy (welshfella):

now multiply first equation by -3 and add -36a - 6b = -294 84a + 6b = 726 adding:- 48a = 432 a = 432/48 = 9 a = 9

OpenStudy (welshfella):

I've found a now you can find b and c gotta go